OpenStudy (anonymous):

HELP r = -6 sin(theta) Part I: Multiply both sides of this equation by r and use the fact that to rewrite this equation in terms of x and y. Part II: Complete the square to produce the final equation. Part III: Using only your answer to part II, what shape does this graph make?

OpenStudy (experimentx):

\[ \sin \theta = \frac y r \text{ and } r = \sqrt{x^2 + y^2 }\] put those values and simplify them, you will get an circle

OpenStudy (experimentx):

\[ x^2 + (y^2 + 6y + 9) - 9 \] also make square by adding 9, other 9 goes to left, giving you radius of 9

OpenStudy (experimentx):

*right

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=polar+plot+r%3D-6sin \theta

OpenStudy (anonymous):

why do u add 9?

OpenStudy (experimentx):

add +9 ... also add -9 to balance them,

OpenStudy (experimentx):

so that you can make \[ x^2 + (y+3)^2 - 9 = 0\]

OpenStudy (anonymous):

sso that would be the answer for tpart b?

OpenStudy (anonymous):

i still a little unclear on part a

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

what part a and b?? i see i ii iii

OpenStudy (anonymous):

im sorry i meaqnt i and ii

OpenStudy (experimentx):

1) just substitute those values of sin \theta (first) then multiply by r on both sides, then put the values of r .. check out my first post for that

OpenStudy (experimentx):

2) from 1) you will be getting \( x^2 + y^2 + 6y = 0 \) completing squrares meand making of this form \( x^2 + (y+3)^2 - 9 = 0 \) which i did earlier ... 3) or you have \( \( x^2 + (y+3)^2 = 3^2 \) and if you remember corrrectly the standard equationf o circle, it's a circle with center 0, -3 and radius 3

OpenStudy (anonymous):

would x be -6 and y be 1?

OpenStudy (experimentx):

?? x is never -6

OpenStudy (experimentx):

plugin those values to see if it satisfies that equation or not